# Optimal Control of a Subdiffusion model using Dirichlet-Neumann and Neumann-Neumann Waveform Relaxation Algorithms

### v1

@@ 4 Auxiliary Results
Now we need some auxiliary result which are necessary for the followings convergent result of DNWR and NNWR algorithm.
Lemma 4.1 (Theorem 2.10 [5]).
Riemann-Liouville time derivative $D^{\alpha}_{RL}y(t)$ with homogeneous initial condition $I^{1-\alpha}_{RL}y(0^{+})=0$ is equivalent to Caputo fractional time derivative $D^{\alpha}_{C}y(t)$ with zero initial condition $y(0)=0$.
Proof 4.2.
@@
After integrating both side with order $\alpha$ i.e. $I^{\alpha}$ we have $y(t)-y(0)=y(t)-\frac{t^{-\alpha}}{\Gamma(1-\alpha)}I^{(1-\alpha)}y(0)$. Thus $I^{(1-\alpha)}y(0)=0$ implies $y(0)=0$. Inversely from the same relation $y(0)=0$ implies $D^{\alpha}_{C}y(t)=D^{\alpha}_{RL}y(t)$. Hence equivalent.
Lemma 4.3.
Stability of an elliptic pde depends on the absolute minimum eigenvalue of the discretize spatial matrix.
Proof 4.4.
For sake of simplicity consider the 2D Poisson equation $u_{xx}+u_{yy}=f(x,y)$. Any numerical scheme of this pde will be stable if the truncation error coming from the Taylor series expansion of the derivative does not effect the solution error too much. That is if we write in terms of error equation using the central difference scheme, we have:
$\frac{e_{i+1,j}-2e_{i,j}+e_{i-1,j}}{\Delta x^{2}}+\frac{e_{i,j+1}-2e_{i,j}+e_{i,j-1}}{\Delta y^{2}}=\eta_{i,j}.$
Therefore in matrix form we have $Ae=\eta$, where $e$ and $\eta$ are column vectors. Hence $e=A^{-1}\eta$. Therefore the stability depends on the absolute minimum eigenvalue of matrix $A$.
Lemma 4.5.
Eigenvalues of matrix $M=\begin{bmatrix}A&\frac{1}{\sigma}B\\ -B&A\end{bmatrix}$ are the collection of eigenvalues of $A+\frac{i}{\sqrt{\sigma}}B$ and $A-\frac{i}{\sqrt{\sigma}}B$, where $A$ and $B$ are block matrix of order $n$.
Proof 4.6.
We know
$\begin{vmatrix}A&\frac{1}{\sigma}B\\ -B&A\end{vmatrix}=\frac{1}{\sqrt{\sigma^{n}}}\begin{vmatrix}\sqrt{\sigma}A&\frac{1}{\sqrt{\sigma}}B\\ -B&A\end{vmatrix}=\frac{1}{\sqrt{\sigma^{n}}}\begin{vmatrix}\sqrt{\sigma}A+iB&\frac{1}{\sqrt{\sigma}}B-iA\\ -B&A\end{vmatrix}$
Then using the determinant formula when $A$ is invertiable, we have
$\begin{vmatrix}A&B\\ C&D\end{vmatrix}=det(A)det(D-CA^{-1}B)$
we have
$\frac{1}{\sqrt{\sigma^{n}}}\begin{vmatrix}\sqrt{\sigma}L+iB&\frac{1}{\sqrt{\sigma}}B-iL\\ -B&L\end{vmatrix}$ $=\frac{1}{\sqrt{\sigma^{n}}}det\left(\sqrt{\sigma}L+iB\right)det\left(L+B(\sqrt{\sigma}L+iB)^{-1}\left(\frac{1}{\sqrt{\sigma}}B-iL\right)\right)$
$=det\left(L+\frac{i}{\sqrt{\sigma}}B\right)det\left(L-\frac{i}{\sqrt{\sigma}}B\right).$
Therefore, when $A$ is invertiable we have
$det\left(\begin{bmatrix}A&\frac{1}{\sigma}B\\ -B&A\end{bmatrix}\right)=det\left(A+\frac{i}{\sqrt{\sigma}}B\right)det\left(A-\frac{i}{\sqrt{\sigma}}B\right)$
But this formula does not contain any $A$ inverse, so we could extend it when $A$ is not invertiable. For this we need to consider the polynomial $P(x)$ of degree $2n$ defined by
$P(x):=det\left(\begin{bmatrix}A-xI_{n}&\frac{1}{\sigma}B\\ -B&A-xI_{n}\end{bmatrix}\right)-det\left(A-xI_{n}+\frac{i}{\sqrt{\sigma}}B\right)det\left(A-xI_{n}-\frac{i}{\sqrt{\sigma}}B\right)$
which has infinitely many zeros when $x$ is not the eigenvalue of $M$, therefore $P(x)$ is a zero polynomial. Hence
$det\left(\begin{bmatrix}A-xI_{n}&\frac{1}{\sigma}B\\ -B&A-xI_{n}\end{bmatrix}\right)=det\left(A-xI_{n}+\frac{i}{\sqrt{\sigma}}B\right)det\left(A-xI_{n}-\frac{i}{\sqrt{\sigma}}B\right).$
Which implies the eigenvalues of matrix $M$ are the collection of eigenvalues of $A+\frac{i}{\sqrt{\sigma}}B$ and $A-\frac{i}{\sqrt{\sigma}}B$.
From the above result we know that the matrix $M$ has the same set of eigenvalues as of $A+\frac{i}{\sqrt{\sigma}}B$ and $A-\frac{i}{\sqrt{\sigma}}B$. Now we want to know is there any connection among them or not. The following result gives the in depth understanding about that.
Lemma 4.7.
@@
Now if we choose $X_{11}=I,X_{12}=I,X_{21}=i\sqrt{\sigma}I,X_{22}=-i\sqrt{\sigma}I$, where $I$ is the identity matrix of order $n$ then left and right side will be equal. Hence $X=\begin{bmatrix}I&I\\ i\sqrt{\sigma}I&-i\sqrt{\sigma}I\end{bmatrix}$ which is invertiable so the matrix $M$ is similar to matrix $\bar{M}$ by the matrix by $X$.
The above result conclude that the diagonalization of matrix $M$ is separately depends on the diagonalization of matrix block $A+\frac{i}{\sqrt{\sigma}}B$ and $A-\frac{i}{\sqrt{\sigma}}B$. The next result give the lower bound on the real part of eigenvalues.
Lemma 4.9.
@@
Proof 4.10.
To prove this result we use the general version of Lyapunov’s theorem on stability [6], which state that: Let $X$ be a complex matrix of order $n$ and $Y$ be a positive definite Hermitian matrix of order $n$. Then there exist a negative definite Hermitian matrix $Z$ for which $XZ+ZX^{*}=Y$ (${}^{\prime}*^{\prime}$ imply the conjugate transpose) holds, if and only if all the eigenvalues of $X$ have negative real part.
Now if we choose $Z=-I$ i.e. negative identity matrix and $X=-M$ then we have $Y=(A+iB)+(A^{T}-iB^{T})=A+A^{T}$ as $B=B^{T}$. From assumption $(A+A^{T})$ is positive definite so the real part of the eigenvalues of $-M$ are negative. Hence the real part of the eigenvalues of $M$ are strictly positive.
Corollary 4.11.
Let $A$ is the first order backward difference matrix with step size $\Delta t$ for $n$ nodes and $B$ is the unit anti-diagonal matrix of same order, i.e.
@@
Clearly $B$ is symmetric and $-(A+A^{T})$ is the central difference matrix with the eigenvalues $\lambda_{j}=-\frac{4}{\Delta t^{2}}\sin^{2}(\frac{\pi}{2}\frac{j}{n+1}),$ $j=1,2,\cdots n$. Therefore the eigenvalues of $A+A^{T}$ are positive, hence the matrix is positive definite. Therefore the real part of the eigenvalues of $M=A+iB$ are strictly positive.
Remark 4.12.
[…]
Corollary 4.13.
Let $D^{\alpha}_{apx}u^{m}$ be the L1 approximation of Caputo derivative of $D^{\alpha}_{t}u(t_{m})$ on the temporal graded mesh $0=t_{0}<t_{1}<\cdots<t_{n}=T$, [8]. Then
[…]
$=\frac{1}{\Gamma(2-\alpha)}\sum_{j=0}^{m-1}\frac{u^{j+1}-u^{j}}{t_{j+1}-t_{j}}\left[(t_{m}-t_{j})^{1-\alpha}-(t_{m}-t_{j+1})^{1-\alpha}\right]$
$=\frac{1}{\Gamma(2-\alpha)}\left[d_{m,1}u^{m}-d_{m,m}u^{0}+\sum_{j=1}^{m-1}\left(d_{m,j+1}-d_{m,j}\right)u^{m-j}\right]$
@@
[…]
Now $(t_{m}-s)^{-\alpha}$ is monotonic increasing function for $s<t_{m}$ and $10<\alpha<$ 1. Hence by using integral mean value theorem one get $0<d_{m,j+1}<d_{m,j}$ for all $m$ and $0\leq j\leq m$. Therefore the difference matrix corresponding to the L1 scheme takes the form
$A=\begin{bmatrix}d_{1,1}&&&&\\ d_{2,2}-d_{2,1}&d_{2,1}&&&\\ \vdots&\ddots&\ddots&&\\ d_{n-1,n-1}-d_{n-1,n-2}&\cdots&&d_{n-1,1}&\\ d_{n,n}-d_{n,n-1}&\cdots&&d_{n,2}-d_{n,1}&d_{n,1}\end{bmatrix}.$
Clearly the $m^{th}$ row sum of absolute elements of matrix $A$ except the diagonal element is $R_{m}:=|(d_{m,2}-d_{m,1})|+\cdots+|(d_{m,m}-d_{m,m-1})|=d_{m,1}-d_{m,m}<d_{m,1}$. Hence by using Gershgorin circle theorem we can say that the eigenvalues of $A+A^{T}$ is strictly positive. Therefore using lemma 4.9 we can conclude that the real part of the eigenvalues of matrix $M=A+iB$, where $B$ takes the form as in corollary 4.11, are strictly positive.
Lemma 4.14.
Let $M=Q+P$ , where $M,Q,P$ are square matrices of order $n$ in $\mathbb{C}$. And $Q$ is diagonalizable, i.e. $Q=XDX^{-1}$. Then every eigenvalue of $M$ lies within the radius $R_{i}$, where $R_{i}$ is the Gershgorin radius of $R=X^{-1}PX$, centered at the addition of some eigenvalue of $Q$ and corresponding diagonal element of $R$.
Proof 4.15.
We have $M=Q+P$ and Q is diagonalizable. Therefore
$M$ $=XDX^{-1}+P$
$X^{-1}MX$ $=D+X^{-1}PX$
$=D+R.$
We know eigenvalues of $M$ and eigenvalues of $X^{-1}MX$ are same. Now we use Gershgorin circle theorem idea to prove the result. Let $\lambda$ be the eigenvalue of $X^{-1}MX$ associate with the eigenvector $y=(y_{j})$ and let $y_{i}$ is the largest absolute value of $y$. Then we have
$X^{-1}MXy$ $=\lambda y$
$(D+R)y$ $=\lambda y.$
Considering the $i^{th}$ equation from the system using $D=(d_{ij})_{n\times n}$ and $R=(r_{ij})_{n\times n}$ matrix elements we have
$\sum_{j=1}^{n}(d_{ij}+r_{ij})y_{j}$ $=\lambda y_{i}$
$d_{ii}y_{i}+\sum_{j=1}^{n}r_{ij}y_{j}$ $=\lambda y_{i}$
$\sum_{j=1,j\neq i}^{n}r_{ij}y_{j}$ $=(\lambda-d_{ii}-r_{ii})y_{i}$
Now using the triangular inequality we have
$|\lambda-d_{ii}-r_{ii}|\leq\sum_{j=1,j\neq i}^{n}|r_{ij}|\frac{|y_{j}|}{|y_{i}|}$
As $y_{i}$ is the largest absolute value in $y$ so
$|\lambda-d_{ii}-r_{ii}|$ $\leq\sum_{j=1,j\neq i}^{n}|r_{ij}|=R_{i}.$
@@
Lemma 4.16.
Let $M=S+J$, where $S$ and $J$ are respectively the lower triangular matrix and unit anti-diagonal matrix of order $n$. Then Greshgorin radius can be easily found out from the relation (14) for $n$ even and from the relation (15) for $n$ odd.
Proof 4.17.
[…]
@@
Corollary 4.18.
Let $S$ be the first order backward difference matrix so takes the same form as in 4.11. For $n$ even case: $A_{m,m}=C_{m,m}=I_{m,m}-L_{m,m}$, where $L_{m,m}$ is lower shift matrix. So,

### v2

@@ 5 Auxiliary Results
Now we need some auxiliary result which are necessary for the convergent of DNWR and NNWR algorithm.
Lemma 1 (Theorem 2.10 [mophou2011optimal]).
The Riemann-Liouville time derivative $D^{\alpha}_{RL}y(t)$ with homogeneous initial condition $I^{1-\alpha}_{RL}y(0^{+})=0$ is equivalent to the Caputo fractional time derivative $_{0}D_{t}^{\alpha}y(t)$ with zero initial condition $y(0)=0$.
Proof.
@@
After integrating both side with order $\alpha$ i.e. $I^{\alpha}$ we have $y(t)-y(0)=y(t)-\frac{t^{-\alpha}}{\Gamma(1-\alpha)}I^{(1-\alpha)}y(0)$. Thus $I^{(1-\alpha)}y(0)=0$ implies $y(0)=0$. Inversely from the same relation $y(0)=0$ implies $D^{\alpha}_{C}y(t)=D^{\alpha}_{RL}y(t)$. Hence equivalent. ∎
Lemma 2.
Stability of an elliptic pde depends on the least absolute eigenvalue of the discretize spatial matrix.
Proof.
The proof is straight forward, hence omitted. ∎
Lemma 3.
@@
Now if we choose $X_{11}=I,X_{12}=I,X_{21}=i\sqrt{\sigma}I,X_{22}=-i\sqrt{\sigma}I$, where $I$ is the identity matrix of order $n$ then left and right side will be equal. Hence $X=\begin{bmatrix}I&I\\ i\sqrt{\sigma}I&-i\sqrt{\sigma}I\end{bmatrix}$ which is invertiable so the matrix $M$ is similar to matrix $\bar{M}$ by the matrix by $X$. ∎
The above result conclude that the diagonalization of matrix $M$ is separately depends on the diagonalization of matrix block $A+\frac{i}{\sqrt{\sigma}}B$ and $A-\frac{i}{\sqrt{\sigma}}B$. The next result gives the lower bound on the real part of eigenvalues.
Lemma 4.
@@
Proof.
For this result we apply the general version of Lyapunov’s theorem on stability [ostrowski1962some], which state that: Let $X$ be a complex matrix of order $n$ and $Y$ be a positive definite Hermitian matrix of order $n$. Then there exist a negative definite Hermitian matrix $Z$ for which $XZ+ZX^{*}=Y$ ($^{\prime}*^{\prime}$ imply the conjugate transpose) holds, if and only if real part of all the eigenvalues of $X$ are negative.
Now if we choose Z=IZ=-I i.e. negative identity matrix and X=MX=-M then we have $Y=(A+iB)+(A^{T}-iB^{T})=A+A^{T}$ as $B=B^{T}$. From assumption $(A+A^{T})$ is positive definite so the real part of the eigenvalues of M-M are negative. Hence the real part of the eigenvalues of $M$ are strictly positive. ∎
Corollary 5.
Let $A$ is the first order backward difference matrix with step size $\Delta t$ for $n$ nodes and $B$ be the exchange matrix of same order, i.e.
@@
Clearly $B$ be symmetric and $(A+A^{T})$ be the central difference matrix with the eigenvalues $\lambda_{j}=-\frac{4}{\Delta t^{2}}\sin^{2}(\frac{\pi}{2}\frac{j}{n+1}),$ $j=1,2,\cdots n$. Therefore the eigenvalues of $A+A^{T}$ are positive, hence the matrix is positive definite. Therefore the real part of the eigenvalues of $M=A+iB$ are strictly positive.
Corollary 6.
Let $D^{\alpha}_{apx}u^{m}$ be the L1 approximation of Caputo derivative of $D^{\alpha}_{t}u(t_{m})$ on the temporal graded mesh $0=t_{0}<t_{1}<\cdots<t_{n}=T$, [stynes2021survey]. Then
$D^{\alpha}_{apx}u^{m}$ $=\frac{1}{\Gamma(1-\alpha)}\sum_{j=0}^{m-1}\int_{t_{j}}^{t_{j+1}}\frac{du}{ds}(t_{m}-s)^{-\alpha}ds\hskip 28.45274pt\textit{ for }m=1,2,\cdots,M$
$=\frac{1}{\Gamma(2-\alpha)}\sum_{j=0}^{m-1}\frac{u^{j+1}-u^{j}}{t_{j+1}-t_{j}}\left[(t_{m}-t_{j})^{1-\alpha}-(t_{m}-t_{j+1})^{1-\alpha}\right].$
@@
(13) $d_{m,j}:=\frac{(t_{m}-t_{m-j})^{1-\alpha}-(t_{m}-t_{m-j+1})^{1-\alpha}}{t_{m-j+1}-t_{m-j}}=\frac{1-\alpha}{t_{m-j+1}-t_{m-j}}\int_{t_{m-j}}^{t_{m-j+1}}(t_{m}-s)^{-\alpha}ds,$
we have
$D^{\alpha}_{apx}u^{m}=\frac{1}{\Gamma(2-\alpha)}\left[d_{m,1}u^{m}-d_{m,m}u^{0}+\sum_{j=1}^{m-1}\left(d_{m,j+1}-d_{m,j}\right)u^{m-j}\right].$
Now $(t_{m}-s)^{-\alpha}$ is monotonic increasing function for $s<t_{m}$ and $0<\alpha<1$. Hence by using integral mean value theorem one get $0<d_{m,j+1}<d_{m,j}$ for all $m$ and $0\leq j\leq m$. Therefore the difference matrix corresponding to the L1 scheme takes the form
(14) $A=\begin{bmatrix}d_{1,1}&&&&\\ d_{2,2}-d_{2,1}&d_{2,1}&&&\\ \vdots&\ddots&\ddots&&\\ d_{n-1,n-1}-d_{n-1,n-2}&\cdots&&d_{n-1,1}&\\ d_{n,n}-d_{n,n-1}&\cdots&&d_{n,2}-d_{n,1}&d_{n,1}\end{bmatrix}$.
Clearly the $m^{th}$ row sum of absolute elements of matrix $A$ except the diagonal element be $R_{m}:=|(d_{m,2}-d_{m,1})|+\cdots+|(d_{m,m}-d_{m,m-1})|=d_{m,1}-d_{m,m}<d_{m,1}$. Hence by using Gershgorin disk theorem we can say that $A+A^{T}$ be the strictly positive definite matrix. Therefore using lemma 4 we can conclude that the real part of the eigenvalues of matrix $M=A+iB$, where $B$ takes the form as in corollary 5, are strictly positive.
Lemma 7.
Let $M=A+B$ , where $M,A,B$ are square matrices of order $n$ in $\mathbb{C}$. Let $B$ is diagonalizable, i.e. $B=XDX^{-1}$. Then each eigenvalue of matrix $M$ is contained within at least one of the Gershgorin discs of matrix $X^{-1}MX$.
Proof.
We have $M=A+B$ and $B$ is diagonalizable by $X$. Therefore
$M$ $=A+XDX^{-1}$
$X^{-1}MX$ $=X^{-1}AX+D$
$=:R+D.$
As $M$ and $X^{-1}MX$ are similar matrices. Hence using Gershgorin circle theorem we conclude that each eigenvalue of matrix $M$ is contained within at least one of the Gershgorin discs $\mathcal{D}(r_{ii}+d_{ii},R_{i})$, where $D=(d_{ij})_{n\times n}$, $R=(r_{ij})_{n\times n}$ and $R_{i}=\sum_{j=1,j\neq i}^{n}|r_{ij}|$. ∎
@@
Lemma 8.
Let $M=S+J$, where $S$ and $J$ are respectively the lower triangular matrix and exchange matrix of order $n$. Then Greshgorin radius can be easily found out from the relation (15) for $n$ even and from the relation (16) for $n$ odd.
Proof.
Let $n$ be even and define $m:=\frac{n}{2}$. Lower triangular matrix $S=\begin{bmatrix}A_{m,m}&O_{m,m}\\ B_{m,m}&C_{m,m}\end{bmatrix}$. $O_{m,m}$ be the zero matrix. $J$ be the exchange matrix which is one form of permutation matrix therefore it can be written in the cycle notation as $(1\,n)(2\,(n-1))...(\lceil\frac{n}{2}\rceil\,\lceil\frac{n+1}{2}\rceil)$. When $n$ is even there are $\frac{n}{2}$number of 2-cycle. Let $(\alpha\,\beta)$ is one of such 2-cycle. Then this 2-cycle contributes to the spectrum the eigenvectors $\tilde{x}_{1}$ and $\tilde{x}_{2}$ corresponding to the eigenvalues $1$ and 1-1 respectively, where both the eigenvectors $\tilde{x}_{1}$ and $\tilde{x}_{2}$ have only $2$ nonzero components at $\alpha$ and $\beta$ positions with value $\frac{1}{\sqrt{2}},\,\frac{1}{\sqrt{2}}$ and $\frac{1}{\sqrt{2}},\,-\frac{1}{\sqrt{2}}$, i.e. $\tilde{x}_{1}=\frac{1}{\sqrt{2}}[0\,\cdots 0\,1_{\alpha^{th}}\,0\cdots\,0\,1_{\beta^{th}}\,0\cdots\,0]$ and $\tilde{x}_{2}=\frac{1}{\sqrt{2}}[0\,\cdots 0\,1_{\alpha^{th}}\,0\cdots\,0\,-1_{\beta^{th}}\,0\cdots\,0]$. Therefore the eigenvector matrix of $J$ can be written as $X=\frac{1}{\sqrt{2}}\begin{bmatrix}I_{m,m}&J_{m,m}\\ J_{m,m}&-I_{m,m}\end{bmatrix}$ corresponding to the eigenvalue matrix $\Lambda_{J}=\begin{bmatrix}I_{m,m}&O_{m,m}\\ O_{m,m}&-I_{m,m}\end{bmatrix}$. As $J$ is also symmetric matrix so it is diagonalizable by orthogonal matrix. Hence $J=X\Lambda_{J}X^{T}$. Therefore using the lemma 7 we can find the Gershgorin radius from the matrix
@@
Corollary 9.
Let $S$ be the Caputo finite difference matrix as in (14) and $n$ is odd and $m:=\frac{n-1}{2}$. From relation (16) we have: